Calculate the oxidation number of chromium in an unknown salt ,if the electrolysis of the salt produce 9.71g of chromium at the cathode after the current of 10A was passed through the cell for 1.5hrs (Cu=52g/mol)


1
Expert's answer
2021-04-21T06:09:37-0400

Solution:

Current, I = 10 A

Time, t = 1.5 hrs = 90 min = 5400 s

Quantity of electricity, Q = I × t = (10 A) × (5400 s) = 54000 C

Mass of chromium (Cr), m = 9.71 g

Molar mass of chromium (Cr), M = 52 g/mol


Faraday's 1st law: The mass of a substance produced at an electrode during electrolysis is proportional to the number of moles of electrons (the quantity of electricity) transferred at that electrode.

Faraday's 1st law can be expressed as:

m = (Q × M) / (Z × F)

where:

m = mass of material (g)

M = molar mass of material (g/mol)

Q = quantity of electricity (C)

Z = number of electrons

F = Faraday constant, 96485 (C mol-1)


Thus:

Z = (Q × M) / (m × F)

Z = (54000 C × 52 g/mol) / (9.71 g × 96485 C mol-1) = 2.997 = 3

Z = 3


Answer: The oxidation number of chromium (Cr) n an unknown salt is +3.